OpenStudy (e.cociuba):

Which is more concentrated: a 1.0 L solution with 20g of sucrose, C12 H22 O11, or a 1.0 L solution with 20 g of glucose, C6 H12 O6? Explain. (It has some math to it, and I no it's not supposed to be in this category, but can someone plz help?)

OpenStudy (e.cociuba):

@electrokid

OpenStudy (anonymous):

this is "chemistry" please post them there. you can still invite me to have fun with the problems and I can try to help out :)

OpenStudy (anonymous):

you can still tag people from the other place ..not necessarily in Math section

OpenStudy (anonymous):

ok, about the problem, did you calculate the molarities?

OpenStudy (anonymous):

@e.cociuba

OpenStudy (e.cociuba):

U just told me to post it on chem....

OpenStudy (zpupster):

sucrose = C(12x12) + H(22 x1) + O(11x16) = 342 20 g/342g/mol= .058mol M or molarity or concentration = .058moles/L do the same with glucose and what ever number is greater is of a greater concentratration

OpenStudy (e.cociuba):

k thanks :)) @zpupster